Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 615     Entry time: 26 May 2025, 09:56
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 26.05.2025 
General Announcement:
- A. Galler mentions the V0 was operated by another group after being interlock on Saturday by a vacuum issue in XTD6 while VAC-OCD was investigating the issue. 
Operation by anyone else has to be prevented in future and only VAC staff shall operate it.

CTRL:
- Issue at MID GUI server being investigated.

DA:
- Nothing major to report.

EEE:
- Nothing to report.

VAC:
- Issue with SA2 Saturday handled by VAC-OCD and PRC. 
- Several mails came from Karabo issue with vacuum pumps in SA1 - to be forwarded to CTRL and discussed.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Not present.

XRO:
- Nothing major to be reported.

LAS:
- Not present.

UNSYS:
- Nothing to report.

SPB:
- Continue with measurements; some difficulties with sample synchronization; nothing else to report.

FXE:
- Nothing to report.

MID:
- Nothing to report apart of the mentioned vacuum issue.

HED:
- Nothing to report.

SCS:
- Nothing to report.

SQS:
- Not present.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00